# Preprint Notice

### v3

@@ III-A The Latency Killer Paradigm
A prevalent belief in data-intensive computing is that network latency is the ultimate data movement killer. While this thinking is common among experienced IT professionals, our empirical testing and architectural analysis demonstrate that this paradigm requires re-examination. The perceived severity of latency is fundamentally governed by TCP's Bandwidth-Delay Product (BDP), which dictates the congestion window required for full bandwidth utilization.
@@
This paradigm can present significant challenges for distributed, data-intensive engineering and scientific endeavors, an impact particularly relevant in the biopharma and life sciences sector—the top global data generator, with data volume projected to grow at a Compound Annual Growth Rate (CAGR) of 36% through 2025 [44]. The challenges of moving data at this scale are a primary concern for industry leaders, including organizations such as tier-1 biopharma businesses and enterprise computing providers like HPE. Inefficient data movement, if not properly addressed, can negatively impact drug discovery, precision medicine development, and research timelines. However, as we demonstrate, these challenges can be significantly mitigated through architectural and software approaches that reduce sensitivity to latency.
@@
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base [8] have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section III-C. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table II for Ethernet adapter ring buffer settings and Table III for tuned Linux kernel parameters.
@@
The test utilized iperf3, a benchmark tool maintained by the U.S. DOE ESnet [12]. Beginning with version 3.16 (released November 30, 2023 [16]), the tool was re-architected to be genuinely multithreaded, making it significantly easier to tune for high-latency, high-speed links (e.g., ~160 Gbps memory-to-memory over a 200 Gbps ESnet testbed data path) [9].
@@
Section III-C provides descriptions of the testbed design. Among many uses of the testbed, it enables the testing of different latency values automatically (Fig. 1). Such a testbed was first constructed in the former Intel Swindon lab, in Swindon, United Kingdom (U.K.). Note that round-trip time (RTT), as reported by tools like ping [36], is approximately 2 × one-way latency.
@@
For reproducibility, the employed HPE server-based appliance, including network interface card (NIC) settings, tuned Linux kernel parameters, and simulated latency values are described in Tables II-IV. Related reference materials are publicly available in a GitHub repository [19].
@@
Figure 2: iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. 7) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.
@@
TABLE I: Key functional capabilities and core features of the Zettar zx data movement software.
@@
Cluster Architecture with Linear Scaling
@@
Embeddable on NVIDIA and Intel DPUs
Built-in Quality of Service (QoS) Support
Integration with typical IAM E.g., Active Directory (AD). Also, OpenID Connect (OIDC) support
TABLE II: rx and tx ring buffer values for Intel® Ethernet Network Adapter E810-2CQDA2.
@@
Figure 3: Bill of Materials (BOM) and component details for the Core (HPE DL380 Gen 11) and Mini (Minisforum MS A2) appliances, demonstrating the vendor- and form-factor-agnostic unified data movement appliance design which achieves consistent performance across implementations.
TABLE III: Tuned Linux kernel parameters. This specific set was empirically determined to be optimal for the zx data mover’s concurrency model, whereas other data movers may require more extensive tuning. For example, see ESnet fasterdata, Linux Tuning [13].
@@
TABLE IV: Some common latency ranges. 10 ms, 50 ms, and 100 ms were used for latency simulation.
@@
Metropolitan Distance (same city or nearby cities) 1 – 10
Interstate Distance (within the same country) 10 – 50
@@
Resources such as the ESnet fasterdata knowledge base provide invaluable guidance. However, practitioners sometimes apply published tuning parameters as static prescriptions rather than adaptable guidelines. This approach can lead to suboptimal performance when it doesn't account for interactions between specific data mover software, hardware stacks, and application workloads. Effective high-performance data movement requires understanding and adaptation beyond copy and paste. Note that the results in Fig. 1 and the modest parameter set in Tables II-III demonstrate how an architectural approach also can reduce complexity and potential for misconfiguration.
III-B Universal Packet Loss and TCP CCAs
This section provides a systematic examination of two interconnected paradigms to end-to-end data movement: the assumption of universal Packet Loss and TCP CCAs. We begin by quantifying the data volumes supported by common transfer rates, then introduce a conceptual model—the “Drainage Basin Pattern”—that frames the entire spectrum of data movement workflows. Subsequently, we present empirical evidence that challenges the conventional view regarding packet loss and the criticality of TCP CCA selection in high-speed environments. A real-world case study, drawn from a Wall Street Journal report, illustrates the practical consequences of these paradigms. To contextualize the scale of these workflows, Table V quantifies the daily data volume achievable at common network speeds. Finally, we discuss why conflating different classes of workflows often leads to suboptimal outcomes.
TABLE V: Daily data volume achievable at three common network speeds.
Data rate (Gbps) Data moved (TB)/day Notes
@@ III-B Universal Packet Loss and TCP CCAs
10 100
100 1000 aka 1 PetaByte
The end-to-end movement of data is generally perceived as a transparent operation by application users and many stakeholders. Nevertheless, the reality is not so simple. The “Drainage Basin Pattern” illustrates the entire spectrum of data movement workflows.
While most users experience only the “source of the river,” where consumer-grade networks may indeed have packet loss, the “high-speed backbone”—the deep main river channel—is an entirely different environment. This recognition is based on extensive experience since 2015 with 100 Gbps and faster connections, gained through collaboration with the U.S. DOE ESnet and various tier-1 organizations. Consequently, our observations confirm that in the well-engineered 100 Gbps+ Research and Education (R&E) networks discussed here—principally ESnet and Internet2 [31]—packet loss is effectively negligible during normal operation. This observed reliability is also a key characteristic of peer R&E networks such as GÉANT [24], AARNet [1], and SingAREN [50].
III-B1 The Packet Loss Paradigm
The nature of data transmission at high speeds differs fundamentally in both engineering and measurement from casual, end-user level data transfers. For high-speed networks, "packets" often vary in size and are transmitted at rates where bit-level impairments dominate over per-packet events. For example, for nodes connected to 10 Gbps and faster networks, it is common to employ "jumbo frame" Maximum Transmission Unit (MTU) of 9,014 or 9,018 bytes, rather than a constant 1,500-byte frame. Consequently, high-speed network device vendors typically specify performance using Bit Error Rates (BER) [39] rather than packet loss. While operators historically reported BERs more explicitly, modern practitioners typically must extract such information indirectly via monitoring infrastructures such as the perfSONAR lookup service [43].
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in [58]) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to 10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
@@ III-B1 The Packet Loss Paradigm
The operational lower bound on error rates is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of [34] (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in [15]. Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This result directly contradicts the notion that packet loss is the dominant limiter in modern, well-engineered wide-area backbones.
III-B2 The TCP CCA Paradigm
@@
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" [6], which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”—a claim that implicitly attributes poor performance to TCP behavior rather than end-system or architectural factors.
@@
It is a telling contrast that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization [15]. This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
@@
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work [15]. Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already documented by RFC8312 [48]. Coincidentally, earlier work [15] described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 [60], titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
@@
It is noteworthy that while Google researchers introduced the model-based BBR CCA in 2018—the same year CUBIC was already the default in widely used Linux operating system distributions such as CentOS 7—BBR has since progressed to version 3 [26]. However, the BBR implementation that ships with RHEL 9.6 remains version 1, which is notably more aggressive than later versions. This distinction is relevant because BBRv1’s aggressiveness should, if anything, favor higher throughput in uncongested environments. Here, we compared the performance of BBR, CUBIC, and Reno (a classical TCP CCA algorithm that emerged in the 1990s) and provided the results in Figs 4-6. The following definitions and methodology are established before presenting the three figures.
@@
Note that statistical analysis for each complete sweep (bulk and streaming) included the calculation of the mean, median, and standard deviation. The mean values are plotted automatically using a custom gnuplot command file [25]. See the GitHub repo [19] for tables containing raw mean, median, and standard deviation values.
Figure 4: A bulk transfer sweep leveraging kTLS (kernel TLS) offload in RHEL 9.6 [47] to evaluate the three default congestion control algorithms (CCAs). BBRv1 did not demonstrate a clear performance benefit over CUBIC or Reno. Incidentally, kTLS did not yield any transfer rate improvement in this case.
Based on the demonstrated functional equivalence of high-speed performance, we conclude that the choice among competing high-speed CCAs is not a primary concern for high-volume, high-throughput data transfer. CUBIC is empirically validated, using the appliances in Fig. 2, as a safe and effective default (Figs. 5-6).
Figure 5: Subsequent bulk transfer sweeps were performed without kTLS due to the prior degradation. Using the default RHEL 9.6 CCAs, BBRv1 and CUBIC exhibited identical throughput across 4 MiB–1 TiB. Reno showed degradation consistent with older congestion control paradigms.
Figure 6: As a unified software data mover [61], zx has built-in streaming capability, motivated by [53]. BBRv1 shows no advantage over CUBIC, even for streaming transfers. kTLS was not used.
@@
III-B3 Practical Implications and Real-World Consequences
The engineering realities discussed thus far have significant, though often invisible, operational and economic impacts. The tangible consequences of suboptimal data transfer are starkly illustrated by recent extreme measures reported by the Wall Street Journal (WSJ) [59]. Given the exponential growth in scientific data, the mobility of multiple petabytes (PBs) is now a common requirement.
A June 12th article [59] titled “Chinese AI Companies Dodge U.S. Chip Curbs by Flying Suitcases of Hard Drives Abroad” detailed how four engineers transported 4.8 PB of training data by flying sixty hard drives from Beijing, China to Kuala Lumpur, Malaysia. This volume of data, sufficient for training large language models (LLMs) such as OpenAI’s ChatGPT, highlights the severe real-world constraints imposed by inefficient electronic data transfer.
A related article from Tom's Hardware [37] clarified that this physical transport was a “meticulously planned operation and took several months of preparation. In sharp contrast, the performance results illustrated in Figs. 4-6, combined with the entries in Table V, demonstrate that transferring this volume of data electronically, using a 100 Gbps link, would require at most one week of sustained transfer.
This dichotomy recalls the hierarchy of solutions presented in Figs. 2-3, whose importance can be illustrated by a simple metaphor: providing water. Watering plants on a balcony requires a watering can. Watering a garden demands a hose. However, supplying a metropolitan area from a reservoir necessitates a massively engineered system, e.g., the Hetch Hetchy Aqueduct [4].

### v4

@@ 3.1. The Latency Killer Paradigm
A prevalent belief in data-intensive computing is that network latency is the ultimate data movement killer. While this thinking is common among many experienced IT professionals, our empirical testing and architectural analysis motivates the re-examination of this paradigm. The perceived severity of latency is fundamentally governed by TCP's Bandwidth-Delay Product (BDP), which dictates the congestion window required for full bandwidth utilization.
@@
This paradigm can present significant challenges for distributed, data-intensive engineering and scientific endeavors, an impact particularly relevant in the biopharma and life sciences sector—the top global data generator, with data volume projected to grow at a Compound Annual Growth Rate (CAGR) of 36% through 2025 (RBC Capital Markets, 2025). The challenges of moving data at this scale are a primary concern for industry leaders, including organizations such as tier-1 biopharma businesses and enterprise computing providers like HPE. Inefficient data movement, if not properly addressed, can negatively impact drug discovery, precision medicine development, and research timelines. However, as we demonstrate, these challenges can be significantly mitigated through architectural and software approaches that reduce sensitivity to latency.
@@
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base (Energy Sciences Network (2025b), ESnet) have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section 3.3. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table 2 for Ethernet adapter ring buffer settings and Table 3 for tuned Linux kernel parameters.
@@
The test utilized iperf3, a benchmark tool maintained by the U.S. DOE ESnet (Energy Sciences Network (2025g), ESnet). Beginning with version 3.16 (released November 30, 2023 (Energy Sciences Network (2023), ESnet)), the tool was re-architected to be genuinely multithreaded, making it significantly easier to tune for high-latency, high-speed links (e.g., ~160 Gbps memory-to-memory over a 200 Gbps ESnet testbed data path) (Energy Sciences Network (2025d), ESnet).
@@
Section 3.3 provides descriptions of the testbed design. Among many uses of the testbed, it enables the testing of different latency values automatically (Fig. 1). Such a testbed was first constructed in the former Intel Swindon lab, in Swindon, United Kingdom (U.K.). Note that round-trip time (RTT), as reported by tools like ping (man7.org, 2025), is approximately 2 × one-way latency.
@@
For reproducibility, the employed HPE server-based appliance, including network interface card (NIC) settings, tuned Linux kernel parameters, and simulated latency values are described in Tables 2-4. More related materials are publicly available in a GitHub repository (Fang, 2025).
@@
Figure 2. iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. 7) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.
A line graph showing iperf3 latency sweep results comparing default vs. tuned kernel settings on a 100 Gbps connection with simulated latencies from 10ms to 100ms.
@@
Table 1. Key functional capabilities and core features of the Zettar zx data movement software.
@@
Cluster Architecture Linear scaling capabilities
@@
Embeddable on specialized hardware Support for NVIDIA and Intel DPUs
Quality of Service (QoS) Built-in support for traffic prioritization
Integration with typical IAM Active Directory (AD) and OpenID Connect (OIDC) support
Table 2. rx and tx ring buffer values for Intel® Ethernet Network Adapter E810-2CQDA2.
@@
Figure 3. Bill of Materials (BOM) and component details for the Core (HPE DL380 Gen 11) and Mini (Minisforum MS A2) appliances, demonstrating the vendor- and form-factor-agnostic unified data movement appliance design which achieves consistent performance across implementations.
A comparison table showing Bill of Materials and component details for Core (HPE DL380) and Mini (Minisforum) data movement appliances.
Table 3. Tuned Linux kernel parameters. This specific set was empirically determined to be optimal for the zx data mover’s concurrency model, whereas other data movers may require more extensive tuning. For example, see ESnet fasterdata, Linux Tuning (Energy Sciences Network (2025e), ESnet).
@@
Table 4. Some common latency ranges. 10 ms, 50 ms, and 100 ms were used for latency simulation.
@@
Metropolitan (same city or nearby) 1 – 10
Interstate (within the same country) 10 – 50
@@
Resources such as the ESnet fasterdata knowledge base provide invaluable guidance. However, practitioners sometimes apply published tuning parameters as static prescriptions rather than adaptable guidelines. This approach can lead to suboptimal performance when it doesn't account for interactions between specific data mover software, hardware stacks, and application workloads. Effective high-performance data movement requires understanding and adaptation beyond copy and paste. Note that the results in Fig. 2 and the modest parameter set in Tables 2-3 demonstrate how an architectural approach also can reduce complexity and potential for misconfiguration.
3.2. Universal Packet Loss and TCP CCAs
This section provides a systematic examination of two interconnected paradigms of end-to-end data movement: the assumption of universal packet loss and the impact of TCP CCA selection. We begin by quantifying the data volumes supported by common transfer rates, then introduce a conceptual model—the “Drainage Basin Pattern”—that frames the entire spectrum of data movement workflows. Subsequently, we present empirical evidence that challenges the two prior assumptions in high-speed environments. A real-world case study, drawn from a Wall Street Journal report, illustrates the practical consequences of these paradigms. To clarify the scale of these workflows, Table 5 quantifies the daily data volume achievable at common network speeds. Finally, we discuss why conflating different classes of workflows often leads to suboptimal outcomes.
Table 5. Daily data volume achievable at three common network speeds.
Data rate (Gbps) Data (TB)/day Notes
@@ 3.2. Universal Packet Loss and TCP CCAs
10 100 High-performance edge transfer
100 1000 i.e. 1 PetaByte
While the end-to-end movement of data often appears transparent to application users, the underlying mechanisms are complex. The “Drainage Basin Pattern” illustrates this full spectrum of data movement workflows, revealing considerations that may not be immediately apparent.
While most users experience only the “source of the river,” where consumer-grade networks may indeed have packet loss, the “high-speed backbone”—the deep main river channel—is an entirely different environment. This recognition is based on extensive experience since 2015 with 100 Gbps and faster connections, gained through collaboration with the U.S. DOE ESnet and various tier-1 organizations. Consequently, our observations confirm that in the well-engineered 100 Gbps+ Research and Education (R&E) networks discussed here—principally ESnet and Internet2 (Internet2, 2025)—packet loss is effectively negligible during normal operation. This observed reliability is also a key characteristic of peer R&E networks such as GÉANT (GEANT Association, 2025), AARNet (AARNet, 2025), and SingAREN (SingAREN, 2025).
3.2.1. The Packet Loss Paradigm
The nature of data transmission at high speeds differs fundamentally in both engineering and measurement from casual, end-user level data transfers. For high-speed networks, "packets" often vary in size and are transmitted at rates where bit-level impairments dominate over per-packet events. For example, for nodes connected to 10 Gbps and faster networks, it is common to employ "jumbo frame" Maximum Transmission Unit (MTU) of 9,014 or 9,018 bytes, rather than a constant 1,500-byte frame. Consequently, high-speed network device vendors typically specify performance using Bit Error Rates (BER) (National Institute of Standards and Technology, 2025) rather than packet loss. While operators historically reported BERs more explicitly, modern practitioners typically must extract such information indirectly via monitoring infrastructures such as the perfSONAR lookup service (perfSONAR, 2025).
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in (Energy Sciences Network (2009), ESnet)) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to 10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
@@ 3.2.1. The Packet Loss Paradigm
The operational lower bound on error rates is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of (Kissel and Fang, 2020c) (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in (Energy Sciences Network (2018), ESnet). Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This result directly contradicts the notion that packet loss is the dominant limiter in modern, well-engineered wide-area backbones.
3.2.2. The TCP CCA Paradigm
@@
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" (Cardwell et al., 2016), which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”—a claim that implicitly attributes poor performance to TCP behavior rather than end-system or architectural factors.
@@
It is a telling contrast that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization (Energy Sciences Network (2018), ESnet). This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
@@
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work (Energy Sciences Network (2018), ESnet). Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already documented by RFC8312 (Rhee et al., 2018). Coincidentally, earlier work (Energy Sciences Network (2018), ESnet) described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 (Xu et al., 2023), titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
@@
It is noteworthy that while Google researchers introduced the model-based BBR CCA in 2018—the same year CUBIC was already the default in widely used Linux operating system distributions such as CentOS 7—BBR has since progressed to version 3 (Google, 2025). However, the BBR implementation that ships with RHEL 9.6 remains version 1, which is notably more aggressive than later versions. This distinction is relevant because BBRv1’s aggressiveness should, if anything, favor higher throughput in uncongested environments. Here, we compared the performance of BBR, CUBIC, and Reno (a classical TCP CCA algorithm that emerged in the 1990s) and provided the results in Figs 4-6. The following definitions and methodology are established before presenting the three figures.
@@
Note that statistical analysis for each complete sweep (bulk and streaming) included the calculation of the mean, median, and standard deviation. The mean values are plotted automatically using a custom gnuplot command file (Gnuplot Development Team, 2024). See the GitHub repo (Fang, 2025) for tables containing raw mean, median, and standard deviation values.
Figure 4. A bulk transfer sweep leveraging kTLS (kernel TLS) offload in RHEL 9.6 (Red Hat, 2025a) to evaluate the three default congestion control algorithms (CCAs). BBRv1 did not demonstrate a clear performance benefit over CUBIC or Reno. Incidentally, kTLS did not yield any transfer rate improvement in this case.
Performance graph showing bulk transfer rates vs. file sizes for BBR, CUBIC, and Reno congestion control algorithms with kTLS offload.
Based on the demonstrated functional equivalence of high-speed performance, we conclude that the choice among competing CCAs is not a primary concern for high-volume, high-throughput data transfer. CUBIC is empirically validated, using the appliances in Fig. 2, as a safe and effective default (Figs. 5-6).
Figure 5. Subsequent bulk transfer sweeps were performed without kTLS due to the prior degradation. Using the default RHEL 9.6 CCAs, BBRv1 and CUBIC exhibited identical throughput across 4 MiB–1 TiB. Reno showed degradation consistent with older congestion control paradigms.
Performance graph showing bulk transfer rates vs. file sizes for BBR, CUBIC, and Reno congestion control algorithms without kTLS.
Figure 6. As a unified software data mover (Zettar Inc., 2025), zx has built-in streaming capability, motivated by (Thayer and Fang, 2019). BBRv1 shows no advantage over CUBIC, even for streaming transfers. kTLS was not used.
Performance graph showing streaming transfer rates vs. file sizes for BBR, CUBIC, and Reno congestion control algorithms.
@@
3.2.3. Practical Implications and Real-World Consequences
The engineering realities discussed thus far have significant, though often invisible, operational and economic impacts. The tangible consequences of suboptimal data transfer are starkly illustrated by recent extreme measures reported by the Wall Street Journal (WSJ) (WSJ Staff, 2025). Given the exponential data growth worldwide, the mobility of multiple petabytes (PBs) is now a common requirement.
A June 12th article (WSJ Staff, 2025) titled “Chinese AI Companies Dodge U.S. Chip Curbs by Flying Suitcases of Hard Drives Abroad” detailed how four engineers transported 4.8 PB of training data by flying sixty hard drives from Beijing, China to Kuala Lumpur, Malaysia. This volume of data, sufficient for training large language models (LLMs) such as OpenAI’s ChatGPT, highlights the severe real-world constraints imposed by inefficient electronic data transfer.
A related article from Tom's Hardware (Morales, 2025) clarified that this physical transport was a “meticulously planned operation and took several months of preparation. In sharp contrast, the performance results illustrated in Figs. 4-6, combined with the entries in Table 5, show that transferring this volume of data electronically, using a 100 Gbps link, would require at most one week of sustained transfer.
This dichotomy recalls the hierarchy of solutions presented in Figs. 2-3, whose importance can be illustrated by a simple metaphor: providing water. Watering plants on a balcony requires a watering can. Watering a garden demands a hose. However, supplying a metropolitan area from a reservoir necessitates a massively engineered system, e.g., the Hetch Hetchy Aqueduct (Bay Area Water Supply and Conservation Agency (2025), BAWSCA).